# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ 4. Results
4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs?
@@
Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes. Each PEFT method was employed to fine-tune the 9 base models (see Section 3.4). Note that we only applied LoRA and $(IA)^{3}$ to Code-T5 as the current PEFT library does not support fine-tuning encoder-decoder models with prompt tuning and prefix tuning. We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 48GB of GPU memory. Additionally, we reported and analyzed the computing resource consumption of these fine-tuning methods, and the results are shown in Figures 4, 5, and 6. Table 5 and Table 6 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC and CM, respectively.
@@
Table 5. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CC (bold: highest values per model, underline: highest values among all models)
Model Method Precision Recall F1 MCC
Small LMs CodeBERT Full FT 90.68% 50.00% 44.86% 0.00%
Prompt Tuning 72.81% 69.55% 70.92% 42.23%
Prefix Tuning 75.82% 73.23% 74.38% 48.98%
LoRA 80.57% 69.03% 72.49% 48.24%
$(IA)^{3}$ 81.22% 68.86% 72.45% 48.53%
GraphCodeBERT Full FT 70.45% 75.65% 72.25% 45.80%
Prompt Tuning 73.88% 71.85% 72.77% 45.69%
Prefix Tuning 81.05% 73.00% 75.95% 53.44%
LoRA 84.72% 69.67% 73.82% 52.27%
$(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
UnixCoder Full FT 67.67% 60.01% 61.69% 26.60%
Prompt Tuning 73.93% 70.64% 72.03% 44.44%
Prefix Tuning 78.73% 73.53% 75.64% 52.00%
LoRA 82.12% 69.00% 72.74% 49.40%
$(IA)^{3}$ 74.33% 69.26% 71.21% 43.29%
CodeT5 Full FT 76.94% 78.09% 77.49% 55.02%
LoRA 63.08% 67.65% 46.82% 30.38%
$(IA)^{3}$ 62.42% 67.82% 63.11% 29.76%
LLMs DeepSeek-Coder-1.3B Prompt Tuning 75.25% 70.27% 72.23% 45.25%
Prefix Tuning 75.07% 75.68% 75.37% 50.75%
LoRA 78.35% 70.51% 73.28% 48.22%
$(IA)^{3}$ 75.74% 70.66% 72.65% 46.11%
DeepSeek-Coder-6.7B Prompt Tuning 74.96% 68.91% 71.13% 43.46%
Prefix Tuning 79.36% 70.72% 73.70% 49.33%
LoRA 75.55% 71.63% 73.26% 47.02%
$(IA)^{3}$ 75.79% 67.29% 69.99% 42.23%
StarCoderBase-1B Prompt Tuning 78.44% 71.06% 73.74% 48.95%
Prefix Tuning 76.84% 70.95% 73.20% 47.42%
LoRA 83.53% 54.82% 54.21% 25.41%
$(IA)^{3}$ 80.02% 70.31% 73.51% 49.39%
StarCoderBase-3B Prompt Tuning 73.67% 71.78% 72.64% 45.41%
Prefix Tuning 74.18% 70.47% 72.01% 44.49%
LoRA 69.39% 54.89% 54.72% 19.49%
$(IA)^{3}$ 79.58% 69.12% 72.39% 47.57%
StarCoderBase-7B Prompt Tuning 73.31% 68.97% 70.69% 42.05%
Prefix Tuning 74.73% 72.62% 73.58% 47.30%
LoRA 69.67% 65.46% 67.03% 34.88%
$(IA)^{3}$ 75.27% 76.48% 75.84% 51.73%
CodeLlama-7B Prompt Tuning 74.10% 71.92% 72.90% 45.97%
Prefix Tuning 71.88% 71.13% 71.49% 43.00%
LoRA 76.57% 73.20% 74.66% 49.66%
$(IA)^{3}$ 71.26% 71.57% 71.41% 42.83%
@@
Table 6. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CM (bold: highest values per model, underline: highest values among all models)
Model Method Precision Recall F1 MCC
Small LMs CodeBERT Full FT 78.55% 53.84% 48.59% 20.95%
Prompt Tuning 84.14% 88.18% 85.20% 72.21%
Prefix Tuning 85.98% 88.42% 86.95% 74.36%
LoRA 85.07% 86.07% 85.53% 71.13%
$(IA)^{3}$ 86.17% 87.32% 86.70% 73.48%
GraphCodeBERT Full FT 78.34% 82.63% 77.01% 60.81%
Prompt Tuning 84.75% 87.83% 85.84% 72.52%
Prefix Tuning 85.54% 86.79% 86.11% 72.33%
LoRA 84.92% 86.97% 85.77% 71.86%
$(IA)^{3}$ 84.15% 86.40% 85.05% 70.51%
UnixCoder Full FT 76.24% 74.36% 75.13% 50.57%
Prompt Tuning 84.64% 87.26% 85.64% 71.85%
Prefix Tuning 86.89% 88.79% 87.70% 75.66%
LoRA 86.99% 90.66% 88.20% 77.56%
$(IA)^{3}$ 86.26% 87.82% 86.95% 74.07%
CodeT5 Full FT 85.66% 89.44% 86.83% 75.00%
LoRA 78.08% 52.26% 45.44% 15.92%
$(IA)^{3}$ 42.00% 43.21% 42.36% -14.73%
LLMs DeepSeek-Coder-1.3B Prompt Tuning 85.83% 88.05% 86.74% 73.85%
Prefix Tuning 86.40% 89.81% 87.57% 76.13%
LoRA 85.14% 87.79% 86.15% 72.88%
$(IA)^{3}$ 85.23% 87.59% 86.17% 72.78%
DeepSeek-Coder-6.7B Prompt Tuning 86.22% 88.92% 87.26% 75.08%
Prefix Tuning 85.73% 89.96% 86.86% 75.57%
LoRA 84.98% 87.48% 85.95% 72.42%
$(IA)^{3}$ 85.58% 83.14% 84.18% 68.68%
StarCoderBase-1B Prompt Tuning 85.68% 87.74% 86.54% 73.39%
Prefix Tuning 84.82% 87.05% 85.72% 71.83%
LoRA 84.81% 87.57% 85.84% 72.33%
$(IA)^{3}$ 86.30% 89.30% 87.40% 75.53%
StarCoderBase-3B Prompt Tuning 84.53% 86.24% 85.26% 70.74%
Prefix Tuning 85.10% 86.38% 85.68% 71.48%
LoRA 84.83% 89.44% 85.77% 74.13%
$(IA)^{3}$ 86.46% 88.38% 87.28% 74.81%
StarCoderBase-7B Prompt Tuning 83.64% 88.07% 84.53% 71.57%
Prefix Tuning 85.35% 87.26% 86.16% 72.59%
LoRA 85.64% 88.45% 86.70% 74.04%
$(IA)^{3}$ 85.72% 88.38% 86.75% 74.05%
CodeLlama-7B Prompt Tuning 74.05% 87.68% 86.00% 72.60%
Prefix Tuning 84.10% 84.36% 84.23% 68.46%
LoRA 85.39% 88.29% 86.46% 73.62%
$(IA)^{3}$ 84.07% 88.40% 85.05% 72.34%
4.1.1. Small LMs
@@
Generally, GraphCodeBERT is the most effective in identifying CC, UnixCoder excels at detecting CM, and CodeT5 with PEFT methods performs the worst in classifying both types of method-level code smells. For CC and CM, CodeBERT demonstrates comparable results to GraphCodeBERT. For example, the F1-score and MCC for CodeBERT with prompt tuning in detecting CM are 85.20% and 72.21%, respectively, while the F1-score and MCC for GraphCodeBERT with prompt tuning for CM are 85.84% and 72.52%. These results are consistent with previous studies (arshad2022codebert, ; rabbani2022comparative, ; lu2021codexglue, ) that CodeBERT and GraphCodeBERT are powerful in code semantics learning, showing outstanding performance in Java code clone detection. GraphCodeBERT is built upon CodeBERT and additionally considers the inherent structure of code (i.e., data flow) in the pre-training stage (guo2021graphcodebert, ). Therefore, it is reasonable that GraphCodeBERT manifests a bit stronger capability than CodeBERT in code smell detection. UnixCoder also shows similar performance to GraphCodeBERT for CC, while it stands out as the best small model for CM, achieving an F1-score of 88.20% and an MCC of 77.56% with LoRA. For all the small LMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods for both CC and CM, with MCC values well below 50%, and even a negative value when tuned with $(IA)^{3}$ for CM. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection compared to the other models, highlighting the limitations of its PEFT adaptation. In contrast, fully fine-tuned CodeT5 outperforms all other fully fine-tuned small models in detecting both CC and CM. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
@@ 4.1.2. LLMs
DeepSeek-Coder-1.3B stands out as the top-performing LLM, consistently achieving the highest or on-par F1-scores and MCC values across all fine-tuning methods when compared to other LLMs. While StarCoderBase-7B with $(IA)^{3}$ gains higher F1-score and MCC value than DeepSeek-Coder-1.3B in detecting CC, it significantly underperforms when fine-tuned with LoRA. Specifically, the F1-score and MCC for DeepSeek-Coder-1.3B with LoRA are 73.28% and 48.22%, respectively, compared to 67.03% and 34.88% for StarCoderBase-7B. Notably, when fine-tuned with PEFT techniques, models with larger parameter sizes within the same family do not surpass, and in some cases, even slightly underperform their smaller counterparts. For example, when detecting CC, DeepSeekCoder-1.3B gains higher F1-scores and MCC values than DeepSeek-Coder-6.7B across the four PEFT methods, and StarCoderBase-1B with prompt tuning and prefix tuning also outperforms StarCoderBase-7B. This observation aligns with Fan et al.’s research (fan2024exploring, ), which shows that larger models do not always have better performance on code-related tasks.
Figure 4 illustrates the peak GPU memory consumption for the three largest models — DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B — using different PEFT techniques for code smell detection. It is evident that, for both CodeLlama-7B and DeepSeek-Coder-7B, $(IA)^{3}$ requires the highest computational resources, with LoRA ranking second. This observation is partly consistent with Li et al.’s work (li2024comprehensive, ) that $(IA)^{3}$ tends to exhibit the highest memory usage. However, Li et al. (li2024comprehensive, ) also found that LoRA consumes less memory than prefix tuning, which is not the case in our results. Unlike DeepSeek-Coder-6.7B and CodeLlama-7B, StarCoderBase-7B shows that LoRA demands more GPU memory than the other PEFT techniques.
Figure 4. Peak GPU memory of DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B with different PEFT methods on detecting CC and CM
4.1.3. Small LMs vs. LLMs
From Table 5 and Table 6, it can be observed that the best-performing models for detecting CC and CM are GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA, both of which are small LMs. More specifically, small LMs with PEFT techniques, with the exception of CodeT5, generally outperform or achieve comparable performance to LLMs. While Weyssow et al.’s work (weyssow2023exploring, ) highlights that LLMs significantly overtake small LMs in code generation tasks, our findings suggest that, for the binary classification task which identifies method-level code smells, small LMs with fewer number of trainable parameters excel beyond LLMs. This discrepancy might arise because the large capacity of LLMs introduces unnecessary complexity, which can obscure the essential patterns needed for the code smell detection task. In contrast, small LMs appear to possess an optimal capacity for capturing these crucial patterns without being encumbered by excessive complexity (tay2023scaling, ). Another possible explanation for the discrepancy is that small LMs, being encoder-only models, are better suited to classification tasks like code smell detection. Encoder-only models are typically more efficient at tasks that require pattern recognition, such as binary classification, while LLMs are often optimized for generative tasks. This indicates that the advantages of LLMs observed in code generation tasks do not necessarily translate to all types of software engineering tasks, and larger models are not always superior when fine-tuned for different downstream applications.
@@ 4.1.4. Best PEFT Method
Overall, for both small LMs and LLMs in the DeepSeek-Coder family, prefix tuning emerges as the most effective technique for method-level code smell detection among the four PEFT methods investigated in this study. In contrast, for the StarCoderBase family, $(IA)^{3}$ tends to be the most effective technique. Meanwhile, for CodeLlama-7B, LoRA proves to be the best PEFT method. For small models, except for CodeT5, as well as LLMs like DeepSeek-Coder-1.3B and DeepSeek-Coder-6.7B, prefix tuning consistently yields the highest F1-scores and MCC values when detecting CC and CM, demonstrating its generally strong effectiveness in fine-tuning these models for code smell detection. In the StarCoderBase family, LLMs fine-tuned with $(IA)^{3}$ outperform those fine-tuned with other PEFT techniques. For example, when detecting CM, StarCoderBase-1B, StarCoderBase-3B, and StarCoderBase-7B all achieve their highest F1-scores and MCC values using $(IA)^{3}$. However, this does not apply to CodeLlama-7B, where using $(IA)^{3}$ results in the worst performance for detecting CC and the second-worst performance for detecting CM. On the other hand, CodeLlama-7B always achieves its best results on this task with LoRA. These findings contrast with those of previous studies (li2024comprehensive, ; weyssow2023exploring, ), which suggested that a single PEFT method could be the best across all models for a specific downstream task. For method-level code smell detection, the optimal PEFT method varies depending on the model. Notably, prompt tuning is the least effective PEFT method for adapting LMs to detect CC and CM, as most models tuned with prefix tuning show worse performance compared to those tuned with other PEFT techniques.
Figure 5. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CC
Figure 6. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CM
@@
According to Table 5 and Table 6, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for method-level code smell detection. For small LMs, excluding CodeT5, the F1-scores and MCC values of prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ all show varying degrees of improvement over full fine-tuning. Our results align with Liu et al.’s prior study (liu2023empirical, ), confirming that PEFT methods can match or surpass the performance of standard full fine-tuning for code understanding tasks while adjusting fewer number of trainable parameters. One possible reason for this could be that full fine-tuning adjusts all of the model’s parameters, which may lead to overfitting, especially when the training data is limited. In contrast, PEFT methods, by focusing on specific parts of the model, allow for more efficient learning without the risk of overfitting to the task at hand.
PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption. Figure 5 and Figure 6 depict the peak GPU memory usage and corresponding MCC values of fine-tuning the four small models. It can be observed that prefix tuning and prompt tuning require significantly less GPU memory than other fine-tuning methods, yet they can still achieve considerable effectiveness. Notably, prefix tuning attains the highest or second-highest MCC value while using minimal GPU memory, proving that prefix tuning is the most efficient PEFT method for small models on code smell detection. When juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage while achieving better or similar results. This further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Small LMs, with the exception of CodeT5, generally achieve better effectiveness than LLMs when using PEFT methods for code smell detection. The most effective PEFT method varies across different models. All PEFT methods outperform or match the performance of full fine-tuning while requiring less peak GPU memory, highlighting the efficiency of PEFT techniques compared to full fine-tuning for detecting CC and CM.

### v2

@@ 4. Results
4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
4.1.1. Experiment Setup
@@
Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes. Each PEFT method was employed to fine-tune the 9 base models (see Section 3.4). Note that we only applied LoRA and $(IA)^{3}$ to Code-T5 as the current PEFT library does not support fine-tuning encoder-decoder models with prompt tuning and prefix tuning. We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 40GB of GPU memory. Table 6 and Table 7 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC, CM, FE, and DC, respectively.
@@
Table 6. Results of small LMs and LLMs with different PEFT methods on detecting CC and CM (bold: highest values per model, underline: highest values among all models)
Model Method CC CM
F1 MCC F1 MCC
SLMs CodeBERT Full FT 70.54% 44.41% 80.57% 61.19%
Prompt Tuning 62.38% 25.90% 76.13% 54.00%
Prefix Tuning 63.88% 27.87% 81.33% 62.94%
LoRA 73.04% 46.10% 71.00% 49.79%
$(IA)^{3}$ 69.44% 40.49% 81.32% 62.96%
GraphCodeBERT Full FT 74.01% 48.08% 80.66% 61.33%
Prompt Tuning 65.24% 30.81% 78.90% 58.60%
Prefix Tuning 69.82% 39.97% 79.18% 60.46%
LoRA 73.15% 46.30% 80.03% 60.93%
$(IA)^{3}$ 72.50% 45.00% 80.00% 60.45%
UnixCoder Full FT 73.04% 46.10% 82.39% 64.89%
Prompt Tuning 69.35% 38.70% 78.66% 57.95%
Prefix Tuning 71.63% 43.26% 82.71% 65.84%
LoRA 72.19% 46.04% 73.18% 51.52%
$(IA)^{3}$ 73.47% 46.98% 82.23% 64.47%
CodeT5 Full FT 73.01% 46.19% 81.13% 62.74%
LoRA 33.33% 0.00% 83.29% -2.95%
$(IA)^{3}$ 43.38% -4.05% 40.32% 16.70%
LLMs StarCoderBase-1B Prompt Tuning 71.07% 42.21% 80.05% 60.11%
Prefix Tuning 70.10% 40.23% 76.88% 54.01%
LoRA 41.41% -1.62% 45.41% 7.01%
$(IA)^{3}$ 74.27% 49.01% 80.90% 62.01%
StarCoderBase-3B Prompt Tuning 69.12% 38.29% 81.87% 63.85%
Prefix Tuning 67.81% 35.69% 78.25% 56.92%
LoRA 40.13% 7.93% 68.41% 46.00%
$(IA)^{3}$ 74.68% 49.96% 82.07% 64.81%
Deepseek-Coder-1.3B Prompt Tuning 70.21% 40.44% 81.77% 63.75%
Prefix Tuning 69.38% 39.62% 79.06% 58.36%
LoRA 69.44% 40.49% 78.42% 57.96%
$(IA)^{3}$ 72.50% 45.01% 81.87% 63.82%
Deepseek-Coder-6.7B Prompt Tuning 67.93% 35.87% 81.19% 63.11%
Prefix Tuning 68.80% 37.62% 80.09% 60.57%
LoRA 69.98% 40.05% 81.34% 63.63%
$(IA)^{3}$ 70.84% 42.36% 83.34% 66.88%
CodeLlama-7B Prompt Tuning 68.68% 37.92% 81.25% 63.43%
Prefix Tuning 53.41% 15.95% 80.77% 62.08%
LoRA 70.12% 40.70% 80.70% 62.59%
$(IA)^{3}$ 70.48% 42.32% 83.00% 66.17%
@@
Table 7. Results of small LMs and LLMs with different PEFT methods on detecting FE and DC (bold: highest values per model, underline: highest values among all models)
Model Method FE DC
F1 MCC F1 MCC
SLMs CodeBERT Full FT 74.09% 48.80% 81.98% 65.60%
Prompt Tuning 56.17% 13.07% 81.56% 64.57%
Prefix Tuning 62.82% 25.93% 81.81% 65.43%
LoRA 67.44% 36.17% 83.17% 68.48%
$(IA)^{3}$ 66.49% 38.48% 83.30% 68.18%
GraphCodeBERT Full FT 77.40% 54.95% 82.66% 67.42%
Prompt Tuning 66.88% 37.49% 83.75% 68.95%
Prefix Tuning 74.19% 48.39% 89.26% 79.26%
LoRA 72.23% 46.36% 85.85% 72.40%
$(IA)^{3}$ 69.84% 45.48% 85.58% 72.56%
UnixCoder Full FT 77.04% 56.76% 83.51% 69.49%
Prompt Tuning 77.32% 55.30% 83.68% 68.74%
Prefix Tuning 75.80% 51.64% 83.89% 69.28%
LoRA 70.97% 41.94% 83.60% 68.65%
$(IA)^{3}$ 73.95% 49.32% 83.96% 69.44%
CodeT5 Full FT 66.49% 38.48% 83.46% 69.12%
LoRA 54.07% 10.02% 70.03% 49.06%
$(IA)^{3}$ 54.65% 9.76% 46.79% -2.66%
LLMs StarCoderBase-1B Prompt Tuning 72.23% 46.36% 84.66% 71.21%
Prefix Tuning 75.65% 52.30% 85.44% 72.19%
LoRA 78.58% 60.68% 79.30% 58.85%
$(IA)^{3}$ 73.95% 49.32% 84.96% 71.71%
StarCoderBase-3B Prompt Tuning 68.96% 39.74% 83.63% 68.47%
Prefix Tuning 67.61% 35.78% 86.56% 74.06%
LoRA 70.85% 42.29% 83.70% 68.64%
$(IA)^{3}$ 69.35% 38.73% 84.90% 71.46%
DeepSeek-Coder-1.3B Prompt Tuning 78.98% 58.34% 84.17% 69.99%
Prefix Tuning 66.88% 37.49% 82.44% 64.90%
LoRA 78.98% 58.34% 82.72% 67.68%
$(IA)^{3}$ 83.87% 67.74% 85.84% 72.57%
DeepSeek-Coder-6.7B Prompt Tuning 78.89% 58.83% 85.56% 72.72%
Prefix Tuning 80.56% 61.81% 83.03% 68.14%
LoRA 77.32% 55.30% 83.43% 68.64%
$(IA)^{3}$ 77.32% 55.30% 85.34% 72.34%
CodeLlama-7B Prompt Tuning 72.52% 45.37% 85.49% 72.55%
Prefix Tuning 63.57% 30.67% 82.30% 64.61%
LoRA 77.32% 55.30% 83.10% 68.31%
$(IA)^{3}$ 74.19% 48.39% 83.54% 68.45%
4.1.2. SLMs
@@
Overall, GraphCodeBERT proved the most effective in identifying CC and DC, while UnixCoder excelled at detecting CM and FE. In comparison, CodeT5 fine-tuned with PEFT methods performed the worst across all types of code smells. CodeBERT performed similarly to GraphCodeBERT, albeit slightly worse. For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively. This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert, ). Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding. When fine-tuned with PEFT methods, the MCC values of CodeBERT even became negative, reaching -4.05% for LoRA on CC and -2.66% for $(IA)^{3}$ on DC. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection, highlighting the limitations of its PEFT adaptation. In contrast, the fully fine-tuned CodeT5 demonstrated acceptable performance in code smell detection. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
@@ 4.1.3. LLMs
DeepSeek-Coder-1.3B stands out as the top-performing LLM, consistently achieving the highest or on-par F1-scores and MCC values across all fine-tuning methods when compared to other LLMs. Although the StarCoderBase models perform similarly to the DeepSeek-Coder models in detecting FE and DC, which involve class-level code understanding, their results are notably less stable for detecting CC and CM, which require method-level understanding. In particular, the performance of StarCoderBase models fine-tuned with LoRA degrades significantly on the detection of CC and CM. Notably, when fine-tuned with PEFT techniques, models with larger parameter sizes within the same family do not surpass, and in some cases, even slightly underperform their smaller counterparts. For example, in FE detection, StarCoderBase-1B and DeepSeek-Coder-1.3B generally achieve higher F1-scores and MCC values than StarCoderBase-3B and DeepSeek-Coder-6.7B.
4.1.4. SLMs vs. LLMs
For the detection of CC and CM, SLMs and LLMs demonstrate generally comparable performance. Notably, in the case of DC, GraphCodeBERT, a representative SLM, significantly outperforms all other models, including LLMs. These results suggest that smaller models can remain competitive, or even superior, for certain types of code smells. In contrast, for FE, LLMs fine-tuned with PEFT methods substantially outperform SLMs, likely due to the more complex and semantic nature of this smell type, which benefits from the richer contextual representations and broader knowledge captured by larger models.
Figure 5. Peak GPU memory usage for detecting CC: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
Additionally, Figure 5 summarizes the peak GPU memory usage of both SLMs and LLMs with different PEFT methods on CC detection. Comparing Figure 5 (a) and (b), we can observe that PEFT methods impose substantially higher peak GPU memory requirements on LLMs compared to SLMs, particularly for larger-sized LLMs.
4.1.5. Method-level vs. Class-level Code Understanding
The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.3.1). By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding. Notably, the models even achieve better performance on class-level code smells, as evidenced by the fact that the best-performing model yields significantly lower results on CC compared to the other smell types.
@@ 4.1.6. Best PEFT Method
Generally, for SLMs, both $(IA)^{3}$ and prefix tuning demonstrate strong performance across different models. For LLMs, $(IA)^{3}$ consistently delivers the best results on the detection of all four types of code smells. To be specific, the most effective PEFT method for SLMs on the task of code smell detection varies depending on the model: for instance, CodeBERT performs optimally with $(IA)^{3}$, whereas UnixCoder tends to achieve better results with prefix tuning. However, for LLMs, $(IA)^{3}$ always outperforms other PEFT methods across all studied code smells. Whether for StarCoderBase Models, Deepseek-Coder Models, or CodeLlama-7B, they all achieve superior performance when fine-tuned using $(IA)^{3}$. Additionally, for LLMs, prompt tuning can be employed as an alternative option. Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases. An example of this is using LoRA on StarCoderBase-1B for CM detection. A case in point is the use of LoRA on StarCoderBase-1B for detecting CM. While StarCoderBase-1B achieves a minimum MCC of 54.01% with the other three PEFT methods (i.e., prompt tuning, prefix tuning, and $(IA)^{3}$), its performance drops sharply to 7.01% with LoRA, resulting in a disparity of over 40%.
@@
According to Table 6 and Table 7, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection. For small LMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning for most code smell types. However, for FE, all PEFT methods perform significantly worse than full fine-tuning. These results suggest that, in most cases, PEFT methods can match or surpass the performance of standard full fine-tuning for code smell detection while adjusting fewer number of trainable parameters.
PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption. Taking CC detection as an example, Figure 5 (a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs. It can be observed that juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC. Among different PEFT methods, both $(IA)^{3}$ and prefix tuning are most effective for SLMs, while $(IA)^{3}$ consistently yields the best results for LLMs. All PEFT techniques outperform or match the performance of full fine-tuning for several code smell types, while also reducing peak GPU memory usage, demonstrating their efficiency. However, for the detection of FE, full fine-tuning remains more effective.